OpenStudy (anonymous):

simplied form of the expression √1/169 1/338 1/13 13 2/169 helpp!

OpenStudy (anonymous):

\[\sqrt{\frac{1}{169}}\] ?

OpenStudy (anonymous):

The answer is : 

OpenStudy (anonymous):

If so, remember: \[\sqrt{\frac{a}{b}} = \frac{\sqrt(a)}{\sqrt(b)}\]

OpenStudy (anonymous):

13 * 13 = 169

OpenStudy (anonymous):

its square root of 1/169

OpenStudy (anonymous):